# Is Deep Learning finally better than Decision Trees on Tabular Data?

### v2

@@ A.1 CatBoost
learning_rate Float [$10^{-5}$, 1]     ✓
@@ A.2 XGBoost
min_child_weight Float [$10^{-8}$, $10^{5}$]     ✓
@@
learning_rate Float [$10^{-5}$, 1]     ✓
@@
gamma Float [$10^{-8}$, $10^{2}$]     ✓
reg_lambda Float [$10^{-8}$, $10^{2}$]     ✓
reg_alpha Float [$10^{-8}$, $10^{2}$]     ✓
@@
We utilized the official XGBoost implementation [^6] https://xgboost.readthedocs.io/en/stable/. While the data preprocessing steps were consistent across all methods, a notable exception was made for XGBoost. For this method, we implemented one-hot encoding on categorical features, as XGBoost does not inherently process categorical values.
@@
A.3 FT-Transformer
@@
d_ffn_factor Float [$\frac{2}{3}$, $\frac{8}{3}$]
lr Float [$10^{-5}$, $10^{-3}$]     ✓
weight_decay Float [$10^{-6}$, $10^{-3}$]     ✓
@@ A.4 SAINT
lr Float [$10^{-5}$, $10^{-3}$]     ✓
weight_decay Float [$10^{-6}$, $10^{-3}$]     ✓
@@ A.6 ResNet
lr Float [$10^{-5}$, $10^{-2}$]     ✓
weight_decay Float [$10^{-6}$, $10^{-3}$]     ✓
@@ A.7 MLP-PLR
lr Float [$10^{-5}$, $10^{-3}$]     ✓
weight_decay Float [$10^{-6}$, $10^{-3}$]     ✓
@@
A.8 RealMLP
For our RealMLP experiments, we use the official implementation  [^9] https://github.com/dholzmueller/pytabkit. Following the authors’ recommendations, we impute missing values using the mean of the training split before applying their preprocessing pipeline. We adopt the recommended default hyperparameters and search space, detailed in Table 9. Additionally, we extend the search space for the initialization standard deviation of the first embedding layer and tune the embedding dimensions, as done for the MLP, to achieve optimal performance.
@@ A.8 RealMLP
A.9 XTab
For XTab, we utilize the official implementation [^10] https://github.com/BingzhaoZhu/XTab. To ensure comparability with other methods, we decouple XTab from AutoGluon and apply the same preprocessing and training pipeline as used for the other models. The original work reports results for both light finetuning and heavy finetuning, so we introduce this as a categorical hyperparameter. If light_finetuning is set to True, the model is finetuned for only 3 epochs. Otherwise, we follow the same epoch range as for the other methods, i.e., $[10,500]$. Furthermore, we use the checkpoint after $2000$ iterations (iter_2k.ckpt), provided by the authors. Table 10 outlines the complete search space used for XTab during hyperparameter optimization (HPO).
@@ A.9 XTab
lr Float [$10^{-5},10^{-3}$]
weight_decay Float [$10^{-6},10^{-3}$]
@@
epochs Integer $3$ (if light_finetuning=True) or $[10,500]$ (otherwise)
A.10 CARTE
@@
lr Float [$10^{-5},10^{-3}$]
weight_decay Float [$10^{-6},10^{-3}$]
@@ A.11 TP-BERTa
lr Float [$10^{-5},10^{-3}$]
weight_decay Float [$10^{-6},10^{-3}$]
@@
A.12 TabPFN
For TabPFN, we utilized the official implementation from the authors [^13] https://github.com/automl/TabPFN. We followed the settings suggested by the authors and we did not preprocess the numerical features as TabPFN does that natively, we ordinally encoded the categorical features and we used an ensemble size of 32 to achieve peak performance as suggested by the authors.
A.13 AutoGluon
@@
A.14 Preprocessing
Our codebase is heavily based on that of (Gorishniy et al., 2021), and we employ the same preprocessing pipeline. Methods for which we do not apply this preprocessing are those that inherently require a different approach, such as TP-BERTa and CARTE, or those implemented within libraries where modifying the preprocessing pipeline is not trivial, such as AutoGluon and RealMLP. In these cases, we use the preprocessing strategies from the original works.
Regarding batch size, we do not tune it in our experiments due to memory constraints. Instead, we determine batch size heuristically similar to (Chen et al., 2024) based on the number of features in the dataset. While batch sizes may vary across datasets, they remain consistent across methods.
Appendix B Raw results tables
B.1 Results after hyperparameter optimization
Table 13 shows the raw results after HPO for CatBoost and XGBoost.
@@ B.1 Results after hyperparameter optimization
Table 13: Average test ROC-AUC per dataset for XGBoost and CatBoost after hyperparameter optimization across CV folds.
@@
Dataset CatBoost XGBoost adult 0.930747 0.930482 analcatdata_authorship 0.999662 0.999816 analcatdata_dmft 0.579136 0.572150 balance-scale 0.972625 0.991268 bank-marketing 0.938831 0.938384 banknote-authentication 0.999935 0.999935 Bioresponse 0.885502 0.888615 blood-transfusion-service-center 0.754965 0.750671 breast-w 0.989162 0.992112 car 1.000000 0.999902 churn 0.922968 0.914432 climate-model-simulation-crashes 0.951480 0.947000 cmc 0.740149 0.735649 cnae-9 0.996316 0.997454 connect-4 0.921050 0.931952 credit-approval 0.934006 0.934692 credit-g 0.801762 0.798571 cylinder-bands 0.912070 0.928116 diabetes 0.837869 0.835638 dna 0.995028 0.995278 dresses-sales 0.595731 0.622414 electricity 0.980993 0.987790 eucalyptus 0.923334 0.918055 first-order-theorem-proving 0.831775 0.834883 GesturePhaseSegmentationProcessed 0.916674 0.916761 har 0.999941 0.999960 ilpd 0.744702 0.748019 Internet-Advertisements 0.979120 0.982276 isolet 0.999389 0.999488 jm1 0.756611 0.759652 jungle_chess_2pcs_raw_endgame_complete 0.976349 0.974087 kc1 0.825443 0.832007 kc2 0.846802 0.843295 kr-vs-kp 0.999392 0.999796 letter 0.999854 0.999819 madelon 0.937562 0.932249 mfeat-factors 0.998910 0.999004 mfeat-fourier 0.984714 0.983375 mfeat-karhunen 0.999264 0.999211 mfeat-morphological 0.965406 0.963075 mfeat-pixel 0.999422 0.999378 mfeat-zernike 0.977986 0.974231 MiceProtein 1.000000 0.999923 nomao 0.996439 0.996676 numerai28.6 0.529404 0.529457 optdigits 0.999844 0.999855 ozone-level-8hr 0.929094 0.922663 pc1 0.875471 0.863061 pc3 0.851122 0.854543 pc4 0.953309 0.951037 pendigits 0.999752 0.999703 PhishingWebsites 0.996482 0.997425 phoneme 0.968024 0.967421 qsar-biodeg 0.930649 0.934875 satimage 0.991978 0.992114 segment 0.996231 0.996126 semeion 0.998687 0.998272 sick 0.998331 0.997950 spambase 0.989935 0.990726 splice 0.995472 0.995049 steel-plates-fault 0.974350 0.972743 texture 0.999948 0.999940 tic-tac-toe 1.000000 0.999710 vehicle 0.943460 0.942080 vowel 0.999259 0.999428 wall-robot-navigation 0.999990 0.999981 wdbc 0.993813 0.994467 wilt 0.990950 0.992192

### v3

@@ A.2 CatBoost
learning_rate Float $10^{-5}$, 1]     ✓
@@ A.3 XGBoost
min_child_weight Float $10^{-8}$, $10^{5}$]     ✓
@@
learning_rate Float $10^{-5}$, 1]     ✓
@@
gamma Float $10^{-8}$, $10^{2}$]     ✓
reg_lambda Float $10^{-8}$, $10^{2}$]     ✓
reg_alpha Float $10^{-8}$, $10^{2}$]     ✓
@@
We utilized the official XGBoost implementation [^4] https://xgboost.readthedocs.io/en/stable/. While the data preprocessing steps were consistent across all methods, a notable exception was made for XGBoost. For this method, we implemented one-hot encoding on categorical features, as XGBoost does not inherently process categorical values, in line with the implementation from the FT-Transformer repository.
@@
A.4 LightGBM
Table 4: Search space for LightGBM.
Parameter Type Range Log Scale
feature_fraction Float [0.5, 1.0]
lambda_l2 Float {0.0, [0.1, 11.0]}     ✓
learning_rate Float [0.001, 1.0]     ✓
num_leaves Integer [4, 768]
min_sum_hessian_in_leaf float [0.0001, 100]     ✓
bagging_fractions Float [0.5, 1.0]
bagging_fractions Float [0.5, 1.0]
n_estimators Integer [100, 2000]
The hyperparameter search space for LightGBM is shown in Table 4. As with other methods, we adopt the preprocessing pipeline from the FT-Transformer repository.
For the default configuration, we retain all library-defined hyperparameters without modification.
A.5 FT-Transformer
@@
d_ffn_factor Float $\frac{2}{3}$, $\frac{8}{3}$]
lr Float $10^{-5}$, $10^{-3}$]     ✓
weight_decay Float $10^{-6}$, $10^{-3}$]     ✓
@@ A.6 SAINT
lr Float $10^{-5}$, $10^{-3}$]     ✓
weight_decay Float $10^{-6}$, $10^{-3}$]     ✓
@@ A.8 ResNet
lr Float $10^{-5}$, $10^{-2}$]     ✓
weight_decay Float $10^{-6}$, $10^{-3}$]     ✓
@@ A.9 MLP-PLR
lr Float $10^{-5}$, $10^{-3}$]     ✓
weight_decay Float $10^{-6}$, $10^{-3}$]     ✓
@@
A.10 TabM
To run TabM in our experiments, we use the pytabkit implementation [^7] https://github.com/dholzmueller/pytabkit. The hyperparameter search space for TabM, presented in Table 10, is adapted from the original work (Gorishniy et al., 2025).
Table 10: Search space for TabM.
Parameter Type Range Log Scale
n_blocks Integer [1, 5]
d_block Integer [64, 1024]
dropout Float [0, 0.5]
hidden_dropout Float [0, 0.5]
lr Float $10^{-4}$, $5\times 10^{-3}$]     ✓
weight_decay Float $10^{-4}$, $10^{-1}$]     ✓
epochs Integer [10, 500]
A.11 RealMLP
For our RealMLP experiments, we use the official implementation in pytabkit7. Following the authors’ recommendations, we impute missing values using the mean of the training split before applying their preprocessing pipeline. We adopt the recommended default hyperparameters and search space, detailed in Table 11. Additionally, we extend the search space for initializating the standard deviation of the first embedding layer and tune the embedding dimensions, as suggested by the authors.
@@ A.11 RealMLP
A.12 XTab
For XTab, we utilize the official implementation [^8] https://github.com/BingzhaoZhu/XTab. To ensure comparability with other methods, we decouple XTab from AutoGluon and apply the same preprocessing and training pipeline as used for the other models. The original work reports results for both light finetuning and heavy finetuning, so we introduce this as a categorical hyperparameter. If light_finetuning is set to True, the model is finetuned for only 3 epochs. Otherwise, we follow the same epoch range as for the other methods, i.e., $[10,500]$. Furthermore, we use the checkpoint after $2000$ iterations (iter_2k.ckpt), provided by the authors. Table 12 outlines the complete search space used for XTab during hyperparameter optimization.
@@ A.12 XTab
lr Float $10^{-5},10^{-3}$]
weight_decay Float $10^{-6},10^{-3}$]
@@
epochs Integer 33 (if light_finetuning=True) or $[10,500]$ (otherwise)
A.13 CARTE
@@
lr Float $10^{-5},10^{-3}$]
weight_decay Float $10^{-6},10^{-3}$]
@@ A.14 TP-BERTa
lr Float $10^{-5},10^{-3}$]
weight_decay Float $10^{-6},10^{-3}$]
@@
A.15 TabPFN
For TabPFN and TabPFNv2, we utilized the official implementations from the authors [^11] https://github.com/automl/TabPFN. We followed the settings suggested by the authors and we did not preprocess the numerical features as TabPFN does that natively, we ordinally encoded the categorical features and we used an ensemble size of 32 for TabPFN to achieve peak performance as suggested by the authors. For TabPFNv2 we use the default settings and do not change anything.
A.16 TabICL
We use the official implementation of TabICL [^12] https://github.com/soda-inria/tabicl and follow the authors’ instructions without modification. In particular, we employ the default checkpoint tabicl-classifier-v1.1-0506.ckpt.
A.17 AutoGluon
@@
Appendix B Hyperparameter analysis
In this section, we first examine the overall importance of hyperparameters for each method, as shown in Figure 9, which quantifies the contribution of each hyperparameter to model performance. The subsequent figures in this section illustrate the effect of individual hyperparameters on the performance metric. The x-axis represents the hyperparameters, while the y-axis denotes the ROC-AUC performance. We calculate hyperparameter importance using the fANOVA (Hutter et al., 2014) implementation in Optuna (Akiba et al., 2019). According to our analysis, the most important hyperparameter for CatBoost is the learning rate, while for XGBoost, it is the subsample ratio of the training instances. For XTab, the learning rate is also the most important hyperparameter, closely followed by the light_finetune hyperparameter, which is a categorical parameter taking values True or False. When light_finetune is True, we fine-tune XTab for only 3 epochs; when it is False, we use the same range of epochs as for the other methods (10 to 500). Similarly, for the MLP with PLR embeddings, the learning rate proves to be the most influential hyperparameter, whereas for RealMLP, the number of units in the hidden layers. For the remaining dataset-specific neural networks in the deep learning family, as well as for CARTE, the number of training epochs is the most important hyperparameter, indicating that training duration plays a critical role in their performance.
(a) CatBoost
(b) XGBoost
(c) ResNet
(d) FT-Transformer
(e) SAINT
(f) TabNet
(g) XTab
(h) CARTE
(i) TP-BERTa
(j) MLP
(k) RealMLP
(l) TabM
Figure 9: Hyperparameter importance for various methods
B.1 CatBoost
B.2 ResNet
B.3 MLP-PLR
B.4 RealMLP
B.5 TabM
Figure 14: Effect of all the hyperparameters on model performance for TabM. The x-axis represents the hyperparameter values, while the y-axis shows the corresponding performance.
B.6 XGBoost
B.7 FT-Transformer
B.8 SAINT
B.9 TabNet
B.10 XTab
B.11 CARTE
B.12 TP-BERTa
B.13 HPO influence on a per-model level
In our analysis of hyperparameter optimization (HPO) versus default configurations across various machine learning methods, we observed that HPO generally led to improved performance. The analysis is depicted in Figure 22. This improvement is reflected in the average rank reductions for most methods when HPO was applied. For example, XGBoost’s average rank improved significantly from 1.94 in its default configuration to 1.06 with HPO, and XTab showed a similar enhancement, moving from a rank of 1.96 down to 1.04.
These findings are visually represented in the accompanying plot, which illustrates the performance gains achieved through HPO. An exception to the general trend was observed with TP-BERTa, where the default configuration slightly outperformed the HPO version (average ranks of 1.47 and 1.53, respectively). This anomaly can be attributed to the computational demands of TP-BERTa. Due to its large model size, TP-BERTa was unable to complete the full 100 hyperparameter tuning trials within the allotted 23-hour time frame, often finishing only a few trials. Consequently, the HPO process may have converged to a suboptimal configuration that did not surpass the performance of the default settings.
Appendix C Ablating the Choice of Refitting
In this ablation, we explore whether refitting the model on the combined training and validation sets (after hyperparameter optimization) provides any measurable benefit. The standard procedure, as described in Section 3.2, uses a 10-fold nested cross-validation: we split the data into 10 folds, use 9 folds for inner cross-validation and HPO, then identify the best hyperparameter configuration and refit the model on all 9 folds before testing on the remaining fold.
Due to the computational resources required, we restrict this analysis to four methods: CatBoost, XGBoost, FT-Transformer, and TabM. The results of this ablation study are presented below, comparing performance with and without refitting.
Section 5, Research Question 4 (Figure 7) summarizes the main ablation results. Here, we extend that analysis by comparing the refitting and no-refitting variants head-to-head at the dataset level and conducting statistical tests to assess significance.
Figure 23 illustrates the performance difference between CatBoost with and without refitting across all datasets. The results clearly indicate that, with only a few exceptions, the refitted version consistently outperforms its non-refitted counterpart.
A similar pattern is observed in Figure 24 for XGBoost. Likewise, Figure 25 shows that refitting yields superior performance on most datasets for TabM too. In contrast, Figure 26 reveals that a greater number of datasets favor the non-refitted FT-Transformer. Nonetheless, the majority still benefit from refitting overall.
Figure 24: Performance difference between XGBoost with refitting and XGBoost without refitting across Datasets. Positive values indicate an improvement in ROC-AUC when refitting is applied, while negative values indicate a performance drop.
Figure 25: Performance difference between TabM with refitting and TabM without refitting across Datasets. Positive values indicate an improvement in ROC-AUC when refitting is applied, while negative values indicate a performance drop.
Furthermore, we conducted a Wilcoxon signed-rank test to compare the performance of refitting versus no-refitting across multiple datasets for each method. The statistical results are summarized in Table 15.
Table 15: Statistical Comparison of Refit vs. No-Refit Methods
Method Pair #Datasets Avg. Diff Median Diff Wilcoxon Stat p-value CatBoost vs. CatBoost_noRefit 68 0.0079 0.0016 180.0000 1.298511e-09 XGBoost vs. XGBoost_noRefit 68 0.0144 0.0029 152.0000 4.413043e-10 TabM vs. TabM_noRefit 68 0.0056 0.0010 101.0000 3.395174e-10 FT vs. FT_noRefit 68 0.0035 0.0004 843.0000 9.356765e-02
For CatBoost, we observed an average performance improvement of $0.0079$ when refitting, with a median difference of $0.0016$. The Wilcoxon test yielded a test statistic of $180.0000$ and a highly significant p-value of $1.2985\cdot 10^{-9}$. This strongly suggests that refitting leads to a statistically significant and consistent improvement in CatBoost’s performance across datasets. Given the very low p-value $(p<0.001)$, we can confidently reject the null hypothesis that refitting has no effect.
For XGBoost, the average improvement with refitting was $0.0144$, with a median difference of $0.0029$. The Wilcoxon test statistic was $152.0000$, with a highly significant p-value of $4.4130\cdot 10^{-10}$. These results indicate that, similar to CatBoost, refitting yields a consistent and statistically significant improvement in performance for XGBoost.
For TabM, refitting led to an average performance improvement of $0.0056$ with a median difference of $0.0010$. The Wilcoxon test yielded a test statistic of $101.0000$ and a highly significant p-value of $3.3952\cdot 10^{-10}$. This provides strong evidence that refitting substantially and consistently improves TabM’s performance across datasets. Given the extremely low p-value $(p<0.001)$, we can confidently reject the null hypothesis, concluding that refitting has a significant positive impact on TabM.
Additionally, Table 16 presents the raw results of FT-Transformer, CatBoost, XGBoost and TabM, in comparison to their non-refitted counterparts.
@@ Appendix C Ablating the Choice of Refitting
Table 16: Average test ROC-AUC per dataset for FT, CatBoost, TabM and XGBoost using refitting vs. no refitting across CV folds.
@@
Dataset CatBoost CatBoost_norefit XGBoost XGBoost_norefit FT FT_norefit TabM TabM_norefit adult 0.930747 0.924052 0.930482 0.916441 0.914869 0.915875 0.919662 0.917949 analcatdata_authorship 0.999662 0.999470 0.999816 0.999304 0.999985 0.999566 1.000000 1.000000 analcatdata_dmft 0.579136 0.547691 0.572150 0.575579 0.576947 0.579169 0.576017 0.568896 balance-scale 0.972625 0.962132 0.991268 0.889868 0.999735 0.995086 0.998912 0.998805 bank-marketing 0.938831 0.937464 0.938384 0.933394 0.938198 0.937470 0.941872 0.941008 banknote-authentication 0.999935 0.999979 0.999935 0.999849 1.000000 1.000000 1.000000 1.000000 Bioresponse 0.885502 0.872449 0.888615 0.853483 0.820159 N/A 0.876671 0.865593 blood-transfusion-service-center 0.754965 0.749848 0.750671 0.726849 0.745975 0.748119 0.748538 0.746875 breast-w 0.989162 0.992507 0.992112 0.991595 0.989503 0.989074 0.995845 0.995032 car 1.000000 0.998453 0.999902 0.999412 0.999751 0.999969 1.000000 1.000000 churn 0.922968 0.916146 0.914432 0.912343 0.914596 0.915300 0.929636 0.918852 climate-model-simulation-crashes 0.951480 0.944551 0.947000 0.947724 0.934671 0.933561 0.939969 0.940694 cmc 0.740149 0.735398 0.735649 0.721967 0.739402 0.736959 0.743797 0.738024 cnae-9 0.996316 0.994599 0.997454 0.983546 0.994497 0.994377 0.998100 0.997319 connect-4 0.921050 0.913372 0.931952 0.923175 0.901170 0.921978 0.941654 0.938588 credit-approval 0.934006 0.940661 0.934692 0.943379 0.935798 0.944236 0.934458 0.928719 credit-g 0.801762 0.773381 0.798571 0.794000 0.783048 0.777810 0.790905 0.770190 cylinder-bands 0.912070 0.867995 0.928116 0.866898 0.915494 0.826412 0.926477 0.886880 diabetes 0.837869 0.822365 0.835638 0.823473 0.831108 0.823379 0.829801 0.823632 dna 0.995028 0.994658 0.995278 0.994620 0.990937 0.989937 0.994505 0.994692 dresses-sales 0.595731 0.605008 0.622414 0.582184 0.620033 0.610016 0.642200 0.615764 electricity 0.980993 0.937421 0.987790 0.948766 0.963076 0.957884 0.968731 0.964049 eucalyptus 0.923334 0.916719 0.918055 0.902200 0.923933 0.911772 0.931897 0.920520 first-order-theorem-proving 0.831775 0.811589 0.834883 0.810288 0.796707 0.785106 0.818255 0.805200 GesturePhaseSegmentationProcessed 0.916674 0.779683 0.916761 0.774266 0.895166 0.799810 0.933828 0.908776 har 0.999941 0.999887 0.999960 0.999919 0.999685 0.999706 0.999966 0.999927 ilpd 0.744702 0.731536 0.748019 0.718251 0.751488 0.737753 0.744875 0.739189 Internet-Advertisements 0.979120 0.972513 0.982276 0.978762 0.974513 0.985391 0.985640 0.982167 isolet 0.999389 0.999282 0.999488 0.999225 0.998817 0.999282 0.999750 0.999628 jm1 0.756611 0.742362 0.759652 0.734592 0.709321 0.725904 0.751557 0.735722 jungle_chess_2pcs_raw_endgame_complete 0.976349 0.973983 0.974087 0.852209 0.999975 0.999861 0.999985 0.999945 kc1 0.825443 0.814042 0.832007 0.808270 0.783519 0.803310 0.813763 0.799767 kc2 0.846802 0.841593 0.843295 0.839860 0.832014 0.837281 0.833491 0.813933 kr-vs-kp 0.999392 0.999419 0.999796 0.999785 0.999777 0.999173 0.999652 0.999659 letter 0.999854 0.999802 0.999819 0.992828 0.999919 0.999886 0.999943 0.999906 madelon 0.937562 0.929178 0.932249 0.912814 0.747391 0.793476 0.809941 0.758538 mfeat-factors 0.998910 0.997917 0.999004 0.998767 0.999015 0.998560 0.999700 0.999550 mfeat-fourier 0.984714 0.984229 0.983375 0.981292 0.984511 0.982372 0.988497 0.987389 mfeat-karhunen 0.999264 0.998802 0.999211 0.998908 0.998682 0.997649 0.999521 0.999267 mfeat-morphological 0.965406 0.965867 0.963075 0.945833 0.970198 0.967869 0.969433 0.969514 mfeat-pixel 0.999422 0.999183 0.999378 0.998464 0.997451 0.998448 0.999478 0.999292 mfeat-zernike 0.977986 0.977831 0.974231 0.975436 0.983479 0.981858 0.984997 0.984036 MiceProtein 1.000000 0.999991 0.999923 0.999871 0.999973 1.000000 1.000000 0.999963 nomao 0.996439 0.995329 0.996676 0.995051 0.990908 0.992552 0.994828 0.994329 numerai28.6 0.529404 0.529350 0.529457 0.528295 0.530315 0.527963 0.529336 0.529882 optdigits 0.999844 0.999780 0.999855 0.999738 0.999616 0.999487 0.999939 0.999891 ozone-level-8hr 0.929094 0.923125 0.922663 0.918516 0.919484 0.919689 0.930601 0.926141 pc1 0.875471 0.850199 0.863061 0.853272 0.917591 0.840223 0.889312 0.862341 pc3 0.851122 0.833527 0.854543 0.833489 0.828743 0.835171 0.843468 0.839846 pc4 0.953309 0.945471 0.951037 0.903891 0.934944 0.944674 0.952956 0.950541 pendigits 0.999752 0.999728 0.999703 0.999777 0.999703 0.999668 0.999739 0.999756 PhishingWebsites 0.996482 0.995649 0.997425 0.996704 0.996760 0.996105 0.997636 0.996866 phoneme 0.968024 0.958699 0.967421 0.957712 0.965071 0.957862 0.971200 0.962861 qsar-biodeg 0.930649 0.928167 0.934875 0.917479 0.919584 0.914716 0.937730 0.929856 satimage 0.991978 0.990444 0.992114 0.990114 0.993516 0.992003 0.994291 0.992353 segment 0.996231 0.995441 0.996126 0.994624 0.994124 0.993598 0.994943 0.994696 semeion 0.998687 0.997784 0.998272 0.974216 0.995548 0.996208 0.998425 0.997714 sick 0.998331 0.997520 0.997950 0.995587 0.997937 0.997762 0.997317 0.996299 spambase 0.989935 0.988718 0.990726 0.988292 0.985969 0.983881 0.989244 0.988850 splice 0.995472 0.992511 0.995049 0.997548 0.992276 0.995195 0.995054 0.994709 steel-plates-fault 0.974350 0.968766 0.972743 0.954217 0.959182 0.962215 0.971043 0.966795 texture 0.999948 0.999946 0.999940 0.999834 0.999983 0.999973 0.999997 0.999997 tic-tac-toe 1.000000 0.999952 0.999710 0.999567 0.996152 0.996209 1.000000 1.000000 vehicle 0.943460 0.933394 0.942080 0.929008 0.963362 0.940233 0.965156 0.955308 vowel 0.999259 0.998833 0.999428 0.997587 0.999713 0.999198 0.999966 0.999854 wall-robot-navigation 0.999990 0.999910 0.999981 0.999586 0.999900 0.999870 0.999912 0.999873 wdbc 0.993813 0.991693 0.994467 0.993817 0.993967 0.986203 0.996573 0.994058 wilt 0.990950 0.991393 0.992192 0.991602 0.993047 0.992642 0.994857 0.995517
Appendix D Raw results tables
D.1 Results after hyperparameter optimization
Table 17 shows the raw results after HPO for CatBoost, LightGBM and XGBoost.
Table 17: Average test ROC-AUC per dataset for CatBoost, LightGBM and XGBoost after hyperparameter optimization across CV folds.
Dataset CatBoost LightGBM XGBoost adult 0.930747 0.931261 0.930482 analcatdata_authorship 0.999662 0.999986 0.999816 analcatdata_dmft 0.579136 0.573445 0.572150 balance-scale 0.972625 0.976799 0.991268 bank-marketing 0.938831 0.938470 0.938384 banknote-authentication 0.999935 0.999979 0.999935 Bioresponse 0.885502 0.886734 0.888615 blood-transfusion-service-center 0.754965 0.723635 0.750671 breast-w 0.989162 0.989162 0.992112 car 1.000000 0.999703 0.999902 churn 0.922968 0.913858 0.914432 climate-model-simulation-crashes 0.951480 0.950490 0.947000 cmc 0.740149 0.730807 0.735649 cnae-9 0.996316 0.984404 0.997454 connect-4 0.921050 0.932440 0.931952 credit-approval 0.934006 0.930848 0.934692 credit-g 0.801762 0.795774 0.798571 cylinder-bands 0.912070 0.929117 0.928116 diabetes 0.837869 0.827963 0.835638 dna 0.995028 0.994942 0.995278 dresses-sales 0.595731 0.617323 0.622414 electricity 0.980993 0.989807 0.987790 eucalyptus 0.923334 0.912027 0.918055 first-order-theorem-proving 0.831775 0.833949 0.834883 GesturePhaseSegmentationProcessed 0.916674 0.920034 0.916761 har 0.999941 0.999959 0.999960 ilpd 0.744702 0.711289 0.748019 Internet-Advertisements 0.979120 0.980094 0.982276 isolet 0.999389 0.999401 0.999488 jm1 0.756611 0.753206 0.759652 jungle_chess_2pcs_raw_endgame_complete 0.976349 0.977605 0.974087 kc1 0.825443 0.803456 0.832007 kc2 0.846802 0.843645 0.843295 kr-vs-kp 0.999392 0.999755 0.999796 letter 0.999854 0.999825 0.999819 madelon 0.937562 0.924095 0.932249 mfeat-factors 0.998910 0.999125 0.999004 mfeat-fourier 0.984714 0.983817 0.983375 mfeat-karhunen 0.999264 0.998958 0.999211 mfeat-morphological 0.965406 0.961250 0.963075 mfeat-pixel 0.999422 0.999131 0.999378 mfeat-zernike 0.977986 0.973627 0.974231 MiceProtein 1.000000 1.000000 0.999923 nomao 0.996439 0.996835 0.996676 numerai28.6 0.529404 0.529077 0.529457 optdigits 0.999844 0.999818 0.999855 ozone-level-8hr 0.929094 0.923040 0.922663 pc1 0.875471 0.874508 0.863061 pc3 0.851122 0.843094 0.854543 pc4 0.953309 0.946295 0.951037 pendigits 0.999752 0.999735 0.999703 PhishingWebsites 0.996482 0.997542 0.997425 phoneme 0.968024 0.965147 0.967421 qsar-biodeg 0.930649 0.931863 0.934875 satimage 0.991978 0.991309 0.992114 segment 0.996231 0.996233 0.996126 semeion 0.998687 0.997646 0.998272 sick 0.998331 0.998073 0.997950 spambase 0.989935 0.989995 0.990726 splice 0.995472 0.995103 0.995049 steel-plates-fault 0.974350 0.973788 0.972743 texture 0.999948 0.999890 0.999940 tic-tac-toe 1.000000 1.000000 0.999710 vehicle 0.943460 0.935611 0.942080 vowel 0.999259 0.998466 0.999428 wall-robot-navigation 0.999990 0.999968 0.999981 wdbc 0.993813 0.994729 0.994467 wilt 0.990950 0.985504 0.992192